How would I do -6(r-2)=-18+4r?
Found 2 solutions by solver91311, macston:
Answer by solver91311(24713)



That would depend a great deal on what you want to do to it.

John

My calculator said it, I believe it, that settles it

Answer by macston(5194)

.
Depends on what you want to do.
To solve for r:
.
-6(r-2)=-18+4r
-6r+12=-18+4r Add 18 to each side.
-6r+30=4r Add 6r to each side.
30=10r Divide each side by 10.
3=r
.
ANSWER: r=3